OpenStudy (johnw):

I had the function f(x)=arccos(x^2+1) and received the antiderivative -2x/√(1-(x^2+1)^2). Can you help me form a function table for f(x)=arccos(x^2+1) and -2x/√(1-(x^2+1)^2) to show when they`re defined?

OpenStudy (johnw):

I need to prove when they´re defined somehow, but I have no idea. Maybe a function table is the wrong way to go?

OpenStudy (eliesaab):

arocos(u) is defined for u between -1 and 1

OpenStudy (eliesaab):

\[ x^2+1 \ge 1\]

OpenStudy (eliesaab):

So you only can take x=0

OpenStudy (loser66):

\[arccos (x^2+1)= \alpha\\cos\alpha=x^2 +1\] Take derivative both sides \[-sin \alpha \dfrac{d}{dx}\alpha =2x\\\dfrac{d}{dx}\alpha =\dfrac{-2x}{sin\alpha}\] Now, back to the first line, \(\alpha '= (arccos (x^2+1))'\) So, \((arccos(x^2+1))'=\dfrac{-2x}{sin\alpha}\) Now, \(cos^2\alpha +sin^2\alpha =1\\sin\alpha =\sqrt{1-cos^2\alpha}\) Hence \((arccos\alpha)'=\dfrac{-2x}{\sqrt{1-(x^2+1)^2}}\)

OpenStudy (johnw):

Thank you so much for the help!

OpenStudy (eliesaab):

since \[ 0\le \cos(\alpha)=x^2 +1\le 1\implies x=0 \] So the function is constant and its derivative is zero

OpenStudy (eliesaab):

the function f(x)